# Interacting type-II semi-Dirac quasiparticles

### v1

@@ Abstract
Transport coefficients, such as dynamical conductivity, can be sensitive to the geometry of the Fermi surface [16]. In our case the shape evolves from convex to concave (boomerang) from low to high energy (Figure 3), with an expected associated change in the conductivity’s frequency dependence. We expect that the conclusions reached in this work will remain valid for finite, small particle density (i.e. small filling of the upper band). Similarly to graphene the charge is not renormalized, as reflected in the non-singular behavior of the polarization function. Topological characteristics, such as the Chern number and the anomalous Hall conductivity, are not affected by the interactions. Overall the interacting Dirac liquid is anticipated to behave as a Fermi liquid but with the remarkable evolution of the Fermi surface shape as a function of the long-range interaction, implying a density of states evolution from linear to power one-third and corresponding changes in the Landau levels quantization condition. Can these effects be detected in real materials? We would suggest, for example, that the modification of the interaction strength $\alpha$ by using dielectric screening (i.e. $e^{2}\rightarrow e^{2}/\kappa$), could potentially probe the interaction exponent presented in Figures (4), (6).

### v2

@@ Abstract
Transport coefficients, such as dynamical conductivity, can be sensitive to the geometry of the Fermi surface [27]. In our case the shape evolves from convex to concave (boomerang) with increasing Fermi energy (Fig. 3), with an expected associated change in the conductivity’s frequency dependence. Moreover, it is of direct relevance that nonlinear magnetotransport has been observed to be very sensitive to changes in Fermi surface geometry in 2D semimetals such as WTe2 [20]. The change of convexity with Fermi level induces a sign inversion of the non-linear magnetoresistance, and accompanying current density response. In the context of our study the features of the band structure are driven by interactions, and the chemical potential can easily be manipulated by doping; whereas in the magnetotransport experiments the pertinent features are electron-hole pockets, and the Fermi level is controlled via temperature. However, given the common underlying fermiology, it is reasonable to expect that peculiar magnetotransport and other potentially unusual phenomena may be observed in type-II semi-Dirac systems.
We expect that the conclusions reached in this work will remain valid for finite, small particle density (i.e. small filling of the upper band). Similarly to graphene the charge is not renormalized, as reflected in the non-singular behavior of the polarization function. Topological characteristics, such as the Chern number and the anomalous Hall conductivity, are not affected by the interactions. The interacting Dirac liquid exhibits remarkable evolution of the Fermi surface shape as a function of the long-range interaction and Fermi energy. The system hosts hybrid quasiparticles with both Dirac and semi-Dirac qualities as reflected in the density of states evolution from linear to power one-third, with accompanying changes to the scaling of associated observables. Can these effects be detected and exploited in real materials? We would suggest, for example, that the modification of the interaction strength $\alpha$ by using dielectric screening (i.e. $e^{2}\rightarrow e^{2}/\kappa$, where $\kappa$ is the dielectric constant) can control the density of states exponent presented in Figs. (4), (6), thus presenting a pathway to tuning the Fermi level at which the surface changes geometry. Given the striking reversal of the current density response in WTe2 for instance, we expect that such tunability might lend itself well to a variety of applications. Furthermore, our results can explain the small low-energy linear regime in (TiO2)5(VO2)3, and more generally, may be used to determine the effective interaction strength in such systems. The generation of a constant term induced by interactions in the effective Hamiltonian is a characteristic feature of semi-Dirac fermions of all types.